MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO May 5, 2009.
Published byModified over 5 years ago
Presentation on theme: "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO May 5, 2009."— Presentation transcript:
MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO May 5, 2009
April 27, 2009Commissioning Needs-Soler2 Commissioning Needs After successful powering up of decay solenoid on 24/04/09, we had a set-back: —Decay solenoid chiller fan burnt out morning 28/04/09 —Burning smell detected —Linde fridge now turned off —Mike Courthold contacted Linde: fan sent by Linde (expected in UK on 5/05/09 but will arrive RAL later this week) —Local team to install fan once it arrives this week. —Solenoid has been warming up all week Plans are being put together to fit a header tank for the demineralised water for the Linde DS chiller Upstream Decay Solenoid shielding reinstated. Decay Solenoid Area is off-limits until new radiation survey carried out during beam (scheduled for today 11am?) – no beam at moment, probably this Friday
April 27, 2009Commissioning Needs-Soler3 Commissioning Needs Working target – Stands.. —one in R78 for demo —one in ISIS for data Detectors – —Tracker cryostats o On 28/04/09 (afternoon) fuse for tracker readout system chiller blew. o Fortunately, Malcolm present and a replacement fuse was found. o Temperature went up to 45K in a few hours but after fuse replaced went back down to normal (~6K) by 19:30. o Craig to reinstall alarm systems this week. —CKOV o Still need network connection in the DSA for new temp and humidity sensors. o Spare HV power supply that failed (faulty chip) has been taken to Geneva for repair Air conditioning —Ducting for A/C venting progressing well Daresbury Lab team carrying out electrical work on main distribution boards. All power to magnet chillers isolated so cannot turn on magnets until this work finished by the end of this week North baseplate grouted last week. Intermediate side plate due 7th May Dates for cleaning and decorating MICE Hall: 29 May-3 July. Need to protect all equipment with fire-resistant polyethylene.
April 27, 2009Commissioning Needs-Soler4 Issues Documentation inventory/overhaul - ongoing —Instructions for shifters – old documentation needs updating —Safety – need all relevant documents in one obvious place Computing —Need network connections to progress in Hall —Still cannot print on OKI printer in R76 but solution found with Kevin Dunford (cost £150) —Wiring for wireless in R76 installed but still not available. —Need Windows PC in MLCR – still pending —Need connection to GRID from MLCR – pending Need beamline magnets on to test alarm handler —Since chillers disconnected this week, water cannot be restored to magnets until next week —Pierrick Hanlet at RAL from 26 May: will test alarms then. Ray Gamet took over as MOM on 2 May